OpenStudy (anonymous):

If a ^{2}-b ^{2} = 36 and -b = 4, what is a ^{2} + 2ab + b ^{2}?

OpenStudy (anonymous):

@zzr0ck3r why is it\[\sqrt{20}\]? Isn't \[\sqrt{52}\] = \[2\sqrt{13}\]?

OpenStudy (cggurumanjunath):